OpenStudy (anonymous):

determine the equation of the line,in point-slope form,that runs through points (6,5) and (-2,-3)

OpenStudy (aroub):

-3-5/-2-6 -8/-8 1 y=mx+b ,m=1 , (6,5) 5=1(6) +b , 5=6+b , b=-1 , the equation of the line is y=x-1